OpenStudy (anonymous):

the solution of m/2-3=m-6?

OpenStudy (callisto):

m/2-3=m-6 (m/2) -m = -6+3 -m/2 = -3 m=6

OpenStudy (anonymous):

thank you callisto

OpenStudy (callisto):

i guess you mean